# Internal diffraction dynamics of trilobite molecules

### v1

@@ I Introduction
Notably, the low-energy scattering of a Rydberg electron off of a ground-state atom can result in the formation of ultralong-range Rydberg molecules (ULRMs), which are bound scattering states that feature massive internuclear proportions [11, 16, 17]. Three main classes of ULRMs are low-$l$, trilobite, and butterfly molecules [18, 19] classified by the low or high angular momentum $l$ state of the Rydberg electron. In contrast to energetically isolated, quantum defect splitted low-$l$ ULRM, trilobite molecules are formed when the nearly degenerate set of high-$l$ hydrogenic states of a Rydberg electron are hybridized by a scattering neutral atom called perturber. Butterfly molecules on the other hand, are formed as a consequence of $p$-wave shape resonances that exist in the low-energy scattering of an electron with an alkali metal atom [10]. These intriguing systems have undergone numerous theoretical and experimental investigations that explored their electronic and rovibrational spectra [20, 21, 22, 23, 24, 25, 26, 27, 19, 28, 29, 30], their responses to external fields [31, 32, 33, 34, 35], nonadiabatic effects [36, 37], and the complex angular momentum structure they feature [38, 39, 17, 40, 41, 42]. Recently there has emerged an interest in the dynamics of the consituent nuclei under the influence of these potentials [34, 43].
@@
This article is organized as follows. In section II we provide our theoretical framework describing the internal electronic structure of the ULRM (II.1), the potential energy landscape as well as the relevant nuclear dynamics (II.2). In Section III we present our numerical results, with subsection III.1 highlighting the energetical structure of the trilobite molecule, subsection III.2 addressing its vibrational dynamics and III.3 detailing the elementary experimental schemes that could be the basis to observe the aforementioned vibrational dynamics. Lastly, our conclusions and future research outlook are provided in Section IV.
@@ II Theoretical framework and Methodology
We begin by characterizing the electronic interaction between the constituents of our system and model the electronic Hamiltonian for the ULRM. The salient features of the trilobite state are highlighted and their effects on the vibrational dynamics of the ULRM are discussed. We then present the time-dependent Schrödinger equation that governs the internal nuclear dynamics of the ULRM.
@@ II.1 Electronic interaction and structure
We consider a system consisting of two 87Rb atoms, one of which is excited to a high-lying Rydberg state, bound to each other via the low-energy scattering of the corresponding Rydberg electron with the ground-state atom. The triplet s-wave scattering of the Rydberg electron off of the ground-state atom facilitates an attractive interaction between the two Rb atoms, which is effectively modeled using a Fermi pseudo potential [50, 10] given by (in atomic units):
@@
where r and R are respectively the positions of the Rydberg electron and the ground-state atom relative to the positive Rydberg core (as seen in Fig. 1(a)). Here, $a_{s}^{T}(k)=a_{s}^{T}(0)+\pi\alpha k/3$, is the low-energy triplet s-wave scattering length for electron-atom collisions, with the zero-energy scattering length $a_{s}^{T}(0)=-$16.05 [11, 51], the ground-state atom polarizability $\alpha$ = 319.2 [52], and the excited electron’s wave-number $k$, given semi-classically in atomic units as $k^{2}=2/R-1/n^{2}$. The singlet scattering interaction, being weaker and repulsive is safely ignored [11, 51]. The adiabatic electronic Hamiltonian for the ULRM is then given by $H_{\mathrm{el}}=H_{\mathrm{Ryd}}+V_{\mathrm{en}}$, with $H_{\mathrm{Ryd}}$ describing the Rydberg electron in the potential of it’s parent ionic core.
@@
Figure 3: Onset of wavepacket diffraction. Probability density plots $|\chi(R,t)|^{2}$ and $|\phi(P,t)|^{2}$ of the scattered wavepacket in the position (a-b) and momentum (c-d) representation. The initial state is a Gaussian wavepacket of characteristic width 100 a.u., centered around $R_{0}$=2400 (a) and $R_{0}$= 3000 (b) a.u.  with the corresponding momentum representation presented in (c) and (d) respectively. Panels (e-f) illustrate the corresponding Fourier transform of $V_{osc}(R)$ probed by each of the wavepackets.
@@ II.2 Vibrational dynamics
Our wavepacket is initialized as a Gaussian centered at position $R_{0}$ with width $\sigma$ with zero initial momentum, i.e. $\chi(R,t=0)\propto\exp\left[-\left(R-R_{0}\right)^{2}/(4\sigma^{2})\right]$, and the time dependent probability density function $|\chi(R,t)|^{2}$ is determined. An initially displaced wavepacket in the trilobite potential, will oscillate back and forth, returning to its initial position accounting for the anharmonic character of the potential. During this process, it will scatter off the oscillatory portion of the potential, which acts as a diffraction grating, facilitating the back-scattering (and forward- scattering) of a time-evolving wavepacket. Via the Fourier transform, we also analyze the momentum space wavefunction, $\phi(P,t)=\mathcal{F}\{\chi(R,t)\}$, and the oscillatory potential $\mathcal{F}\{V_{osc}(R)\}$ to study and explain such scattering effects in momentum space. In the following we numerically investigate the vibrational dynamics of ULRM in the position and momentum space for different initial state configurations and different principle quantum numbers, using the framework discussed above.
@@ III.1 Potential energy landscape
Additional insight can be gained by analyzing the oscillatory potential in momentum space through the Fourier transform of $V_{osc}(R)$, $\mathcal{F}\{V_{osc}(R)\}$, the results of which are shown in Fig. 2(c). In momentum space representation, the potentials exhibit a clear side-band structure that corresponds to it’s undulating nature in $R$. For example, the $n$=55 trilobite state exhibits oscillations in $R$ whose wavelength is approximately 100 a.u., though it does vary increasing up to $\sim 230$ a.u in the $R$-range considered here. This results in a Fourier spectrum that is strongly peaked at $P$ = 0.03 a.u., where $P$ is the radial momentum in the Fourier space. As the oscillations of the PEC deviates from a strict periodic behaviour in position space, we observe a broadening of the side-band structures with an overlay of multiple peaks in Fourier space. Note that the Fourier transform is performed using a truncated range of $R$-values ranging from $R$=1300 a.u. to 4500 a.u. This corresponds to the range of internuclear separations probed by the oscillating wave-packet. The extremes of the position of the peak of the wavepacket are determined by the classical turning points of the system in the trilobite potential when the wavepacket is started from rest displaced to one side. The usage of such a limited grid in position space restricts the accuracy of the Fourier analysis and invariably contributes to further broadening of the side-bands and the exact positions of the multiple peaks. Finally, the $n$-dependence of the amplitude of oscillations of the PEC as discussed before, is also reflected in the $n$-dependence of the amplitude and position of the side-bands in Fig. 2(c).
@@ III.2 Vibrational dynamic effects
Figure 4: Probability densities $|\chi(R,t)|^{2}$ and $|\phi(P,t)|^{2}$ of the scattered wavepacket in the position (a) and momentum (b) representation. The initial state is a Gaussian wavepacket of width 100 a.u., centered around $R_{0}$= 4000 a.u. The arrows mark the stickiness of the wavepacket discussed in the text. (c) Time-slices of $|\phi(P,t)|^{2}$, illustrating the back-scattering undergone by the propagating wave-packet.
@@
Onset of wavepacket diffraction - We begin by performing a series of dynamical simulations of the ULRM’s nuclear wavepacket initialized in a Gaussian state centered around different values of $R$, in an $n=55$ trilobite PEC. The wavepackets are constructed such that the width of the corresponding probability density exhibits a standard deviation of 100 a.u., which is comparable in size to the wavelength of the oscillations in the trilobite potential. Figure 3 illustrates the probability density functions of the time evolving wavepacket, in the position space (Fig. 3(a-b)) and momentum space (Fig. 3(c-d)). The position of the initial state determines the total energy of the time evolving wavepacket, thereby putting an upper bound on the magnitude of radial momentum probed by the nuclei during it’s dynamics. A Gaussian state centered around $R_{0}=2400$ a.u (Fig. 3(a,c)), i.e. relatively close to the equilibrium internuclear distance ($R_{eq}\sim$ 2190 a.u), performs low energy oscillations around $R_{eq}$ with complex multi-well interference effects. Interestingly, a further increase in the total energy of the wavepacket, achieved by having it’s initial position extended to $R_{0}=3000$ a.u (Fig. 3(b,d)), exhibits the onset of a clear multiple peak structure that resembles a time propagating diffraction pattern. The emergence of these unique interference patterns is explained by focusing on the momentum picture. Oscillations that achieve lower momentum, are not able to probe the clear side-band structure of the trilobite potential exhibited by it’s oscillatory features (as seen in Fig. 3(e)). Hence they only display very faint forward and back-scattering peaks that are quickly dispersed in the course of time. However, oscillations as portrayed in Fig. 3(d), quite clearly achieve higher momentum beyond the side-band structures (as seen in Fig. 3(f)), thereby showing clear signs of forward and back-scattering in the momentum picture. Such an explanation treats the oscillatory structure of the trilobite potential as a diffraction grating in momentum space.
@@
Internal diffraction dynamics - To emphasize the phenomena discussed above, we analyze the vibrational dynamics of a ULRM with an initial Gaussian wavepacket centered at $R_{0}=4000$ a.u. again with a width of 100 a.u. The resulting time-dependent probability density is shown in Fig. 4(a). In this case the system samples enough of the oscillatory potential to constantly undergo forward and back-scattering of the oscillatory part of the trilobite potential produces a train of roughly evenly spaced side peaks. The origin of this behavior becomes apparent in the momentum space probability density shown in Fig. 4(b). Here, the initial momentum space wavepacket scattering off the oscillatory potential producing adjacent scattered peaks corresponding to the momentum of the peak of the Fourier transform. Figure 4(c) shows time slices of the momentum probability density function at different times showing peaks separated in momentum by around $\sim$ 0.03 a.u, approximately the peak of the side-bands seen in the momentum representation of the PEC in Fig. 2(c). The resulting fragmented wave-train structure in momentum space exhibits a clear and coherent multi-peak structure, reminiscent of that observed in waves scattering from a diffraction grating. Apart from the diffraction effects, we also notice that the probability density plots in figure 4(a) (marked by the arrows), exhibit sustained population near the initial state of the wavepacket which we denote as “stickiness”. This is a result of a large overlap of the initial wavepacket with eigenstates that are localized around a well near the initial position. Part of the wave-packet evolves with the time-scales associated with those localized state, which is larger than the time-scale of anharmonic oscillation, thereby exhibiting said stickiness.
@@
Figure 5: Interference effects in the scattering of a wide wavepacket. Probability densities $|\chi(R,t)|^{2}$ and $|\phi(P,t)|^{2}$ of the scattered wavepacket are shown in the position (a) and momentum (b) representation. The initial state is a Gaussian wavepacket of width 300 a.u., centered around $R_{0}$=4000 a.u. Time-slices of $|\chi(R,t)|^{2}$ and $|\phi(P,t)|^{2}$ illustrating key stages of evolution are shown in the panels (c) and (d), respectively.
III.3 Prototypical experimental setup

### v2

@@ I Introduction
Notably, the low-energy scattering of a Rydberg electron off of a ground-state atom can result in the formation of ultralong-range Rydberg molecules (ULRMs), which are bound scattering states that feature massive internuclear proportions [11, 16, 17]. Three main classes of ULRMs are low-$l$, trilobite, and butterfly molecules [18, 19] classified by the low or high angular momentum $l$ state of the Rydberg electron. In contrast to energetically isolated, quantum defect splitted low-$l$ ULRM, trilobite molecules are formed when the nearly degenerate set of high-$l$ hydrogenic states of a Rydberg electron are hybridized by a scattering neutral atom called perturber. Butterfly molecules on the other hand, are formed as a consequence of $p$-wave shape resonances that exist in the low-energy scattering of an electron with an alkali metal atom [10]. These intriguing systems have undergone numerous theoretical and experimental investigations that explored their electronic and rovibrational spectra [20, 21, 22, 23, 24, 25, 26, 27, 19, 28, 29, 30], their responses to external fields [31, 32, 33, 34, 35], nonadiabatic effects [36, 37, 38], and the complex angular momentum structure they feature [39, 40, 17, 41, 42, 43]. Recently there has emerged an interest in the dynamics of the consituent nuclei under the influence of these potentials [34, 44]. Apart from the neutral trilobite molecule investigated here, nuclear wavepacket dynamics of charged Rydberg molecules have also garnered a great deal of recent interest [45, 46].
@@
This article is organized as follows. In section II we provide our theoretical framework describing the internal electronic structure of the ULRM (II.1), the potential energy landscape as well as the relevant nuclear dynamics (II.2). In Section III we present our numerical results, with subsection III.1 highlighting the energetical structure of the trilobite molecule, subsection III.2 detailing its vibrational dynamics, III.3 addressing potential limitations in the theoretical framework, and III.4 detailing the elementary experimental schemes that could be the basis to observe the aforementioned vibrational dynamics. Lastly, our conclusions and future research outlook are provided in Section IV.
@@ II Theoretical framework and Methodology
We begin by characterizing the electronic interaction between the constituents of our system and model the electronic Hamiltonian for the ULRM. The salient features of the trilobite state are highlighted and their effects on the vibrational dynamics of the ULRM are discussed. We then present the time-dependent Schrödinger equation that governs the internal nuclear dynamics of the ULRM. Throughout our work, the system and the results are presented and analysed in atomic units (a.u.) unless explicitly mentioned otherwise.
@@ II.1 Electronic interaction and structure
We consider a system consisting of two 87Rb atoms, one of which is excited to a high-lying Rydberg state, bound to each other via the low-energy scattering of the corresponding Rydberg electron with the ground-state atom. The triplet s-wave scattering of the Rydberg electron off of the ground-state atom facilitates an attractive interaction between the two Rb atoms, which is effectively modeled using a Fermi pseudo potential [53, 10] given by:
@@
where r and R are respectively the positions of the Rydberg electron and the ground-state atom relative to the positive Rydberg core (as seen in Fig. 1(a)). Here, $a_{s}^{T}(k)=a_{s}^{T}(0)+\pi\alpha k/3$, is the low-energy triplet s-wave scattering length for electron-atom collisions, with the zero-energy scattering length $a_{s}^{T}(0)=-$16.05 [11, 54], the ground-state atom polarizability $\alpha$ = 319.2 [55], and the excited electron’s wave-number $k$, given semi-classically in atomic units as $k^{2}=2/R-1/n^{2}$. The singlet scattering interaction, being weaker and repulsive is safely ignored [11, 54]. The adiabatic electronic Hamiltonian for the ULRM is then given by $H_{\mathrm{el}}=H_{\mathrm{Ryd}}+V_{\mathrm{en}}$, with $H_{\mathrm{Ryd}}$ describing the Rydberg electron in the potential of its parent ionic core.
@@
Figure 3: Onset of wavepacket diffraction. Probability density plots $|\chi(R,t)|^{2}$ and $|\phi(P,t)|^{2}$ of the wavepacket scattering off an $n$=55 trilobite potential, in the position (a-b) and momentum (c-d) representation. The initial state is a Gaussian wavepacket of characteristic width 100 a.u., centered around $R_{0}$=2400 (a) and $R_{0}$= 3000 (b) a.u. with the corresponding momentum representation presented in (c) and (d) respectively. Panels (e-f) illustrate the corresponding Fourier transform of $V_{osc}(R)$ probed by each of the wavepackets.
@@ II.2 Vibrational dynamics
Our wavepacket is initialized as a Gaussian centered at position $R_{0}$ with width $\sigma$ with zero initial momentum, i.e. $\chi(R,t=0)\propto\exp\left[-\left(R-R_{0}\right)^{2}/(4\sigma^{2})\right]$, and the time dependent probability density function $|\chi(R,t)|^{2}$ is determined. An initially displaced wavepacket in the trilobite potential, will oscillate back and forth, returning to its initial position accounting for the anharmonic character of the potential. During this process, it will scatter off the oscillatory portion of the potential, which acts as a diffraction grating, facilitating the back-scattering (and forward- scattering) of a time-evolving wavepacket. Via the Fourier transform, we also analyze the momentum space wavefunction, $\phi(P,t)=\mathcal{F}\{\chi(R,t)\}$, and the oscillatory potential $\mathcal{F}\{V_{osc}(R)\}$ to study and explain such scattering effects in momentum space. In the following we numerically investigate the vibrational dynamics of ULRM in the position and momentum space for different initial state configurations and different principle quantum numbers, using the framework discussed above. The stationary eigenstates of $H_{n}$ are calculated utilizing a tenth order finite difference method [61], which are then used to simulate the time-evolving wavepacket (Eq. 8). Owing to the simplicity of the system, we achieve converged, accurate results with relatively little computational effort. Hence the results have been calculated with step sizes of $\Delta_{R}\sim 2$ a.u. and $\Delta_{t}\sim 0.05$ ns.
@@ III.1 Potential energy landscape
Additional insight can be gained by analyzing the oscillatory potential in momentum space through the Fourier transform of $V_{osc}(R)$, $\mathcal{F}\{V_{osc}(R)\}$, the results of which are shown in Fig. 2(c). In momentum space representation, the potentials exhibit a clear side-band structure that corresponds to its undulating nature in $R$. For example, the $n$=55 trilobite state exhibits oscillations in $R$ whose wavelength is approximately 100 a.u., though it does vary increasing up to $\sim 230$ a.u in the $R$-range considered here. This results in a Fourier spectrum that is strongly peaked at $P$ = 0.03 a.u., where $P$ is the radial momentum in the Fourier space. As the oscillations of the PEC deviates from a strict periodic behaviour in position space, we observe a broadening of the side-band structures with an overlay of multiple peaks in Fourier space. Note that the Fourier transform is performed using a truncated range of $R$-values ranging from $R$=1300 a.u. to 4500 a.u. This corresponds to the range of internuclear separations probed by the oscillating wave-packet. The extremes of the position of the peak of the wavepacket are determined by the classical turning points of the system in the trilobite potential when the wavepacket is started from rest displaced to one side. The usage of such a limited grid in position space restricts the accuracy of the Fourier analysis and invariably contributes to further broadening of the side-bands and the exact positions of the multiple peaks. Finally, the $n$-dependence of the amplitude of oscillations of the PEC as discussed before, is also reflected in the $n$-dependence of the amplitude and position of the side-bands in Fig. 2(c).
@@ III.2 Vibrational dynamic effects
Figure 4: Probability densities $|\chi(R,t)|^{2}$ and $|\phi(P,t)|^{2}$ of the wavepacket scattering off an $n$=55 trilobite potential, in the position (a) and momentum (b) representation. The initial state is a Gaussian wavepacket of width 100 a.u., centered around $R_{0}$= 4000 a.u. The arrows mark the stickiness of the wavepacket discussed in the text. (c) Time-slices of $|\phi(P,t)|^{2}$, illustrating the back-scattering undergone by the propagating wave-packet. (d) The probability densities of the molecular eigenstates, with the localized states responsible for the initial “stickiness” highlighted (purple) near the initial Gaussian (cyan).
@@
Onset of wavepacket diffraction - We begin by performing a series of dynamical simulations of the ULRM’s nuclear wavepacket initialized in a Gaussian state centered around different values of $R$, in an $n=55$ trilobite PEC. The wavepackets are constructed such that the width of the corresponding probability density exhibits a standard deviation of 100 a.u., which is comparable in size to the wavelength of the oscillations in the trilobite potential. Figure 3 illustrates the probability density functions of the time evolving wavepacket, in the position space (Fig. 3(a-b)) and momentum space (Fig. 3(c-d)). The position of the initial state determines the total energy of the time evolving wavepacket, thereby putting an upper bound on the magnitude of radial momentum probed by the nuclei during its dynamics. A Gaussian state centered around $R_{0}=2400$ a.u (Fig. 3(a,c)), i.e. relatively close to the equilibrium internuclear distance ($R_{eq}\sim$ 2190 a.u), performs low energy oscillations around $R_{eq}$ with complex multi-well interference effects. Interestingly, a further increase in the total energy of the wavepacket, achieved by having its initial position extended to $R_{0}=3000$ a.u (Fig. 3(b,d)), exhibits the onset of a clear multiple peak structure that resembles a time propagating diffraction pattern. The emergence of these unique interference patterns is explained by focusing on the momentum picture. Oscillations that achieve lower momentum, are not able to probe the clear side-band structure of the trilobite potential exhibited by its oscillatory features (as seen in Fig. 3(e)). Hence they only display very faint forward and back-scattering peaks that are quickly dispersed in the course of time. However, oscillations as portrayed in Fig. 3(d), quite clearly achieve higher momentum beyond the side-band structures (as seen in Fig. 3(f)), thereby showing clear signs of forward and back-scattering in the momentum picture. Such an explanation treats the oscillatory structure of the trilobite potential as a diffraction grating in momentum space.
@@
Internal diffraction dynamics - To emphasize the phenomena discussed above, we analyze the vibrational dynamics of a ULRM with an initial Gaussian wavepacket centered at $R_{0}=4000$ a.u. again with a width of 100 a.u. As discussed earlier, increasing $R_{0}$ farther away from $R_{eq}$ further increases the momentum probed by the wavepacket allowing it to clearly probe the momentum side-band structure of the potential as it scatters off the PEC. Hence, following the trend from the previous simulations (with $R_{0}=2400,3000$ a.u.), we see that initializing the wavepacket with $R_{0}=4000$ a.u. facilitates repeated coherent back-scattering as the wavepacket traverses through the PEC. The resulting time-dependent probability density is shown in Fig. 4(a). In this case the system samples enough of the oscillatory potential to constantly undergo forward and back-scattering of the oscillatory part of the trilobite potential produces a train of roughly evenly spaced side peaks. The origin of this behavior becomes apparent in the momentum space probability density shown in Fig. 4(b). Here, the initial momentum space wavepacket scattering off the oscillatory potential producing adjacent scattered peaks corresponding to the momentum of the peak of the Fourier transform. Figure 4(c) shows time slices of the momentum probability density function at different times showing peaks separated in momentum by around $\sim$ 0.03 a.u, approximately the peak of the side-bands seen in the momentum representation of the PEC in Fig. 2(c). The resulting fragmented wave-train structure in momentum space exhibits a clear and coherent multi-peak structure, reminiscent of that observed in waves scattering from a diffraction grating. Apart from the diffraction effects, we also notice that the probability density plots in Fig. 4(a) (marked by the arrows), exhibit sustained population near the initial state of the wavepacket which we denote as “stickiness”. This is a result of a large overlap of the initial wavepacket with eigenstates that are localized around a well near the initial position (see Fig. 4(d)). Figure 4(d) shows the vibrational probability densities of eigenstates that participate in the wavepacket dynamics as well the initial wavepacket (cyan) for reference. These localized states (purple) have a large overlap with the initial wavepacket and exhibit a vibrational splitting around $\sim 46$ MHz, as opposed to the delocalised states (grey) exhibiting a vibrational splitting of $\sim 67$ MHz. Part of the wave-packet evolves with the time-scales associated with the localized states ($\sim 210$ ns), which is noticeably larger than the time-scale of the anharmonic oscillation ($\sim 150$ ns), thereby exhibiting said stickiness. Due to the limited amount of eigenstates involved, this particular phenomena continues to occur for even larger values of $t$ ($\geq$ 400 ns), retaining its spatial localisation periodically, similar to a quantum revival. However, the propagating diffraction pattern is continuously scattered off of a band of varying frequencies diminishing the chance of a complete quantum revival in finite time.
@@
Figure 5: Interference effects in the scattering of a wide wavepacket. Probability densities $|\chi(R,t)|^{2}$ and $|\phi(P,t)|^{2}$ of the wavepacket scattering off an $n$=60 trilobite potential, are shown in the position (a) and momentum (b) representation. The initial state is a Gaussian wavepacket of width 300 a.u., centered around $R_{0}$=4000 a.u. Time-slices of $|\chi(R,t)|^{2}$ and $|\phi(P,t)|^{2}$ illustrating key stages of evolution are shown in the panels (c) and (d), respectively.
III.3 Limitations of the theoretical framework
We have ignored the effects of $p$-wave shape resonances and spin-structure in our analysis. Inclusion of the $p$-wave scattering term would result in a more complex PEC structure that features avoided crossings between trilobite and butterfly curves near $R_{eq}$. These avoided crossings would substantially alter the long-time dynamics of the wavepacket as it traverses through the full potential energy landscape. This might lead to the introduction of vibronic decay, surface hopping, diabatic stabilization and several other nonadiabatic effects. Recent works have explored the spectral features of such systems [36, 37, 38] and have shown that nonadiabatic effects are of increasing importance at higher $n$-values (like $n\geq$ 55) [45]. However, our present work is a proof of methodology meant to bring attention and interest to the study of wavepacket dynamics of trilobite molecules. Furthermore, while the $p$-wave shape resonances will substantially alter the long time-scale dynamics, the initial time dynamics should still qualitatively exhibit the diffractive effects discussed. Also, one can mitigate the complexity introduced by spin-structure to some extent by careful preparation and state selection of the hyperfine state of the ground-state atom. Moreover, the hyperfine splitting of the PECs introduced by ground-state 87Rb remains around $\sim$ 6.8 GHz while the depth of the trilobite potentials scale as $n^{-3}$ with increasing $n$. At high enough $n$-values, for example $n\geq$ 55, the rather large energetic separation between the different hyperfine states can be utilised to select and probe a single PEC. Of course, it is important to note that these compilcations can be eliminated by selecting a different ground-state perturbing atom, such as Sr which has no low energy, p-wave electron-atom resonances and lacks any hyperfine structure.
III.4 Prototypical experimental setup